OpenStudy (anonymous):

What is the minimum work needed to push a 1225 kg car 340 m up along a 17.5° incline?

OpenStudy (anonymous):

(a) Ignore friction. ___J (b) Assume the effective coefficient of friction retarding the car is 0.20? ___J

OpenStudy (anonymous):

Turn that incline into a right triangle and then you just have to calculate the vertical height the car actually climbs: sin 17.5 = opposite side / hypotenuse 0.30 = opposite side / 340m opposite side = 0.30 * 340m = 102.2m the opposite side on a right triangle is the y-axis or the vertical height height = 102.2m Work = force * distance work = KE = PE = potential energy of car if it fell from 102.2m down to ground by force of gravity PE = mgh PE = 1225 kg * 9.8 m/s^2 * 102.2m Work = PE = 1,226,911 Joules = 1227 kJ